Foreverangel:

What is the inverse of f(x)=(x−5)^2 for x≥5 where function g is the inverse of function f? g(x)=√x−5, x≥5 g(x)=√x+5, x≥−5 g(x)=√x −5, x≥0 g(x)=√x +5, x≥0

AZ:

To find the inverse of a function we have to switch the x and y and solve for y So let's just change f(x) to y to make things simpler. So, first let's switch x and y y = (x-5)^2 x = (y - 5)^2 All you have to do is swap their places. Now you have to solve for y! So first things first, how will we get rid of the \( ^2\) on the right side? We can take the square root of both sides! What do you get when that happens?